A parallel-plate capacitor has an area of 100 mm2 and the plates are separated by 100 μm . If the capacitor contains paper between the plates, what is its capacitance



1
Expert's answer
2022-05-16T13:40:29-0400
"C=\frac{\kappa\epsilon_0 A}{d}"

"C=\frac{3.7*8.85*10^{-12}*100*10^{-6}}{100*10^{-6}}=3.27*10^{-11}\rm F"